OpenStudy (anonymous):

-5x - 8y = 15, (-3,0) , is this a real solution ? ._.

OpenStudy (anonymous):

yes. Given -5x-8y=15 coordinates (-3,0) Solution: Plug one of the coordinates into the equation. I'm going to use the x value -3 so -5x-8y-15 -5(-3)-8y=15 15-8y=15 Now get all like terms into one side. Remember what you do to one side you must do to the other side. So i'm going to minus 15 from both sides. You'll be left with -8y=0 Now divide y=0 Now do the same thing again but this time plug in 0 for y.